Graph the function
f(x)=-2x-5

Answer by elima(1433)

f(x)=-2x-5
Make a table choosing values for x then solve for y;
let x = -1
f(1)=-2(-1)-5
f(1)=3-5
f(1)=-2
first point (1,-2)
let x = -2
f(-2)=-2(-2)-5
f(-2)=4-5
f(-2)=-1
second point (-2,-1)
let x = 0
f(0)=2(0)-5
f(0)=-5
third point (0,-5)
:)